OpenStudy (anonymous):

The measure of an angle is 40 degrees more than the measure of its supplement. what is the measure of each angle? Please explain

OpenStudy (anonymous):

x + x + 40 = 180 2x + 40 = 180 2x = 140 x = 70 x + 40 = 110